# Generative Medical Event Models Improve with Scale

### v1

@@ Appendix D Event prevalence and co-occurrence
One way to assess the plausibility of CoMET’s generations is to measure the prevalence of generated medical events. Given the model has to learn how to properly generate syntactically valid medical event sequences, it is not necessarily trivial to also learn the frequency at which different events of different types should appear for individual patients. For 20,000 patients, we generated 1 year’s worth of sequences 25 times and measured, on average, how many times different medical events occurred. We measured this for diagnoses (at the 3-character ICD-10-CM code level for simplicity), labs, medications, and procedures, as seen in Figure 20.
In addition to measuring agreement in the prevalence of single events in generated sequences, we next asked whether pairs of events co-occur within a patient’s generated sequence the proper amount. This second-degree question is a quick, aggregate measure that the model understands the relationship between events, both within and across data types. We measured the fraction of patients in which each possible pair of events co-occurred. Results in Figure 20 and Table 6 show that CoMET generally does well at generating events at the same rate of co-occurrence as observed in real patient medical event sequences.
Figure 20: Plausibility of predicted medical event prevalence and pairwise co-occurrence. On the left are heatmap scatter plots for the prevalence of different discrete medical events in CoMET-L’s predicted generations versus the ground truth. The plots include diagnoses (first 3 ICD-10-CM characters only), medications, labs, and procedures. On the right are heatmap scatter plots showing the fraction of patients in which pairs of medical events co-occurred within one year, comparing CoMET-L’s predicted generations versus ground truth. Note that because these are all log-log plots, zero prevalence and co-occurrence values are also shown but with an explicit gap.
@@
Figure 21: T2DM-specific outcome performance measured by AUCROC improves as train loss decreases. We evaluated each CoMET model, along with earlier checkpoints from the CoMET-S and CoMET-M training runs, on all T2DM tasks. We fit a sigmoid curve to all points except for CoMET-L to assess the sigmoid curve’s predictive utility. We evaluated each model using a more conservative $n=20$ simulations. Stars indicate compute-optimal models.
Figure 22: Single-encounter generations improve as train loss decreases. We evaluated each CoMET model, along with earlier checkpoints from the CoMET-S and CoMET-M training runs, on all of the single-encounter completion tasks. We fit a sigmoid curve to all points except for CoMET-L to assess the sigmoid curve’s predictive utility. We evaluated each model using $n=20$ simulations. Stars indicate compute-optimal models.
@@ Appendix F Bias and Fairness
Figure 23: Bias Impact via Subcohort Analysis. CoMET-L’s AUCROCs with 95% bootstrapped CI stratified by demographic groups are displayed for five different predictive tasks for brevity. In order of which they appear they are: one from acute-on-chronic event, one from incident disease risk, two from T2DM-specific outcomes, and one from operational outcomes.
@@ Appendix G Additional Results for Disease-Specific Outcomes
CoMET predicts lab values that exhibit moderate correlation with actual observed values. CoMET predictions are noisy as a consequence of the limited expressiveness of the model’s lab result value bucketed tokens (see Section 5.2.4 for how labs are tokenized).
Figure 24: Heat map scatter plots comparing 3-month CoMET-L predictions of lab values to the patient’s actual results. Plots show HgbA1c for patients with diabetes and show total cholesterol, LDL cholesterol, and HDL cholesterol for patients with hyperlipidemia.
@@
CoMET-L 1.21 1.69 11.57 18.30 31.51 41.91 43.01 58.66
CoMET-M 1.24 1.73 13.15 20.36 33.41 44.75 45.75 61.98
CoMET-S 1.30 1.81 13.62 21.06 33.86 44.70 44.88 60.61
@@
Table 16: MAE and RMSE values for the corresponding subplots in Figure 24 across all CoMET models and the best performing task-specific model.

### v3

@@ Appendix D Event prevalence and co-occurrence
One way to assess the plausibility of Curiosity’s generations is to measure the prevalence of generated medical events. Given the model has to learn how to properly generate syntactically valid medical event sequences, it is not necessarily trivial to also learn the frequency at which different events of different types should appear for individual patients. For 20,000 patients, we generated 1 year’s worth of sequences 25 times and measured, on average, how many times different medical events occurred. We measured this for diagnoses (at the 3-character ICD-10-CM code level for simplicity), labs, medications, and procedures, as seen in Figure 20.
In addition to measuring agreement in the prevalence of single events in generated sequences, we next asked whether pairs of events co-occur within a patient’s generated sequence the proper amount. This second-degree question is a quick, aggregate measure that the model understands the relationship between events, both within and across data types. We measured the fraction of patients in which each possible pair of events co-occurred. Results in Figure 20 and Table 6 show that Curiosity generally does well at generating events at the same rate of co-occurrence as observed in real patient medical event sequences.
Figure 20: Plausibility of predicted medical event prevalence and pairwise co-occurrence. On the left are heatmap scatter plots for the prevalence of different discrete medical events in Curiosity-L’s predicted generations versus the ground truth. The plots include diagnoses (first 3 ICD-10-CM characters only), medications, labs, and procedures. On the right are heatmap scatter plots showing the fraction of patients in which pairs of medical events co-occurred within one year, comparing Curiosity-L’s predicted generations versus ground truth. Note that because these are all log-log plots, zero prevalence and co-occurrence values are also shown but with an explicit gap.
@@
Figure 21: T2DM-specific outcome performance measured by AUCROC improves as train loss decreases. We evaluated each Curiosity model, along with earlier checkpoints from the Curiosity-S and Curiosity-M training runs, on all T2DM tasks. We fit a sigmoid curve to all points except for Curiosity-L to assess the sigmoid curve’s predictive utility. We evaluated each model using a more conservative $n=20$ simulations. Stars indicate compute-optimal models.
Figure 22: Single-encounter generations improve as train loss decreases. We evaluated each Curiosity model, along with earlier checkpoints from the Curiosity-S and Curiosity-M training runs, on all of the single-encounter completion tasks. We fit a sigmoid curve to all points except for Curiosity-L to assess the sigmoid curve’s predictive utility. We evaluated each model using $n=20$ simulations. Stars indicate compute-optimal models.
@@ Appendix F Bias and Fairness
Figure 23: Bias Impact via Subcohort Analysis. Curiosity-L’s AUCROCs with 95% bootstrapped CI stratified by demographic groups are displayed for five different predictive tasks for brevity. In order of which they appear they are: one from acute-on-chronic event, one from incident disease risk, two from T2DM-specific outcomes, and one from operational outcomes.
@@ Appendix G Additional Results for Disease-Specific Outcomes
Curiosity predicts lab values that exhibit moderate correlation with actual observed values. Curiosity predictions are noisy as a consequence of the limited expressiveness of the model’s lab result value bucketed tokens (see Section 5.2.4 for how labs are tokenized).
Figure 24: Heat map scatter plots comparing 3-month Curiosity-L predictions of lab values to the patient’s actual results. Plots show HgbA1c for patients with diabetes and show total cholesterol, LDL cholesterol, and HDL cholesterol for patients with hyperlipidemia.
@@
Curiosity-L 1.21 1.69 11.57 18.30 31.51 41.91 43.01 58.66
Curiosity-M 1.24 1.73 13.15 20.36 33.41 44.75 45.75 61.98
Curiosity-S 1.30 1.81 13.62 21.06 33.86 44.70 44.88 60.61
@@
Table 16: MAE and RMSE values for the corresponding subplots in Figure 24 across all Curiosity models and the best performing task-specific model.